Introduction to Mathematical Modeling Quiz I

Friday, September 7

Problem 1. What is the law of demand? Your answer should include a sentence and a graph.

Problem 2. What is the law of supply? Again, your answer should include a sentence and a graph.

Problem 3. What is meant by an equilibrium solution in supply and demand economics? Label the equilibrium solution on a graph and concisely explain its significance.

Problem 4. If the baseball industry is not operating at an equilibrium point describe what may happen.